OpenStudy (anonymous):

What is the solution of the system? Use a graph. y = x + 5 y = –5x – 1

OpenStudy (lncognlto):

Well, graph em! Where they intersect, that's the solution.

OpenStudy (anonymous):

I don't know how :(

OpenStudy (lncognlto):

Well, then, assign a set of x values for both, say -3 to 3. Make a table for the x values and the y values for each equation once the x has been substituted into it. So for -3, y = x + 5 = -3 + 5 = 2 y = –5x – 1 = -5(-3) -1 = 15 - 1 = 14. Like this: Now complete the table.